Weston Products manufactures an industrial cleaning compound that goes through three processing departments—Grinding, Mixing, and Cooking. All raw materials are introduced at the start of work in the Grinding Department. The Work in Process T-account for the Grinding Department for May is given below:

Work in Process—Grinding Department
Inventory, May 1 21,800 Completed and transferred
to the Mixing Department
?
Materials 133,400
Conversion 225,500
Inventory, May 31 ?

The May 1 work in process inventory consisted of 18,000 pounds with $14,600 in materials cost and $7,200 in conversion cost. The May 1 work in process inventory was 100% complete with respect to materials and 30% complete with respect to conversion. During May, 167,000 pounds were started into production. The May 31 inventory consisted of 15,000 pounds that were 100% complete with respect to materials and 60% complete with respect to conversion. The company uses the weighted-average method in its process costing system.

Required:

1. Compute the Grinding Department's equivalent units of production for materials and conversion in May.

2. Compute the Grinding Department's costs per equivalent unit for materials and conversion for May.

3. Compute the Grinding Department's cost of ending work in process inventory for materials, conversion, and in total for May.

4. Compute the Grinding Department's cost of units transferred out to the Mixing Department for materials, conversion, and in total for May.
Answer #1

Unit completed and transferred out = 18000+ 167000-15000 = 170000

1) Equivalent unit of production

EUP-Material EUP-Conversion
Units transferred out 170000 170000
Ending work in process 15000 15000*60% = 9000
Equivalent unit of production 185000 179000

2) Cost per equivalent unit

Material Conversion
Beginning work in process 14600 7200
Cost added 133400 225500
Total Cost 148000 232700
Equivalent unit 185000 179000
Cost per equivalent unit 0.80 1.30

3) Cost of ending work in process

Material Conversion    total
Cost of ending work in process 15000*0.80 = 12000    9000*1.3 = 11700 23700  

4) Cost of units transferred out

Material Conversion Total
Cost of units transferred out 170000*0.8 = 136000 170000*1.3 = 221000    357000